One solution contains 50% alcohol and another solution contains 80% alcohol. How many liters of each solution should be mixed to make 13.5 liters of a 70% alcohol solution?
Answer by edjones(8007)

Let x be the amount of 50% alcohol required.
.50x + .80(13.5-x)=.70*13.5
.50x+10.8-.80x=9.45
-.3x=-1.35
x=4.5 L
13.5-x=9.0 L
.
Ed